A company sells q ribbon winders per year at $p per ribbon winder. The demand function for ribbon winders is given by p=300−0.02q. Find the elasticity of demand when the price is $70 apiece. Will an increase in price lead to an increase in revenue?


1
Expert's answer
2022-01-12T18:09:06-0500

If P = $70, then:

70 = 300 - 0.02q,

q = 11,500 units.

q = 15,000 - 50p.

The elasticity of demand is:

"Ed = -50*70\/11,500 = 0.304," so the demand is inelastic.

It means that an increase in price will lead to an increase in revenue.